The values g and h are defined as follows: 1+ V5 1-15 g= and h= 2 2 Prove that 1 + g =g2 and 1 + h = h2.
Answer #1

M SOL c of 1+ 5 2 It ght T +T 2 c 2+1ts 2 !! 3 + 5 2 = ge L gr 2 2 (1 + 5) (2)? (15+ (55)?+ 2(1)55 b. = t +5+ 255 4 1) 7 6 +Hence proved 1+ g = g​​​​​​2

h = 1 - 5 2 1th I tI-5 2 2 tl -55 U 2 3 - 5 2 2 2-55 2 2 (1-552 (2) ² (+)² + 1-55² - 2120155 4 ) I + 5 - 255 4 6- 255 4 y (3-

Hence proved 1 + h = h​​​​​​2